1. Methane (CH4) is a highly flammable gas that can combust according to the equation below: 

CH4 (g) + 2 O2 (g) → 2 H2O (g) + CO2 (g)

a.)Given the standard heats of formation (ΔHof) in the table below, calculate the ΔH of the combustion reaction.

Compound ∆Hf°(kJ/mol)

CO2 -393.5

H2O(g) -241.8

H2O (l) -285.8

CH4 -74.87


b.)Use the ΔHcombustion from part (a) and the bond enthalpies in the table below to determine the bond enthalpy per mole of O2


Bond type Average Bond Enthalpy (kJ/mol)

C - H 414

C - O 360

C = O 799

O - H 464

O = O ?


c.)Draw a reaction profile for the combustion of methane. Use a vertical arrow (↕️) to indicate ΔHcombustion that you calculated in (a)







1
Expert's answer
2021-03-08T05:55:47-0500

a) CO2 -393.5

H2O(g) -241.8

H2O (l) -285.8

CH4 -74.87

∆H= -393.5 +(-285.8 x 2) - (-74.87)= 1 039,97 kJ/mol.

b) C - H 414

C - O 360

C = O 799

O - H 464

O = O ?

414 x 4 + 2x = 799 x 2 + 464 x 2

x= 435.

c)